# 3D-Properties: Identifying Challenges in DPO and Charting a Path Forward

### v1

@@ Abstract
Aligning large language models (LLMs) with human preference has recently gained tremendous attention, with the canonical yet costly RLHF-PPO and the simple and straightforward Direct Preference Optimization (DPO) as two examples. Despite the efficiency, DPO has rarely be used in the state-of-the-art production-level LLMs, implying its potential pathologies. In this work, we revisit DPO with a comprehensive examination of its empirical efficacy and a systematic comparison with RLHF-PPO. We identify the 3D-properties of DPO’s learning outcomes: the Drastic drop in the likelihood of rejected responses, the Degradation into LLM unlearning, and the Dispersion effect on unseen responses through experiments with both a carefully designed toy model and practical LLMs on tasks including mathematical problem-solving and instruction following. These findings inherently connect to some observations made by related works and we additionally contribute a plausible theoretical explanation for them. Accordingly, we propose easy regularization methods to mitigate the issues caused by 3D-properties, improving the training stability and final performance of DPO. Our contributions also include an investigation into how the distribution of the paired preference data impacts the effectiveness of DPO. We hope this work could offer research directions to narrow the gap between reward-free preference learning methods and reward-based ones.
@@ 1 Introduction
Large language models (LLMs) trained on extensive datasets have shown outstanding performance across diverse tasks and domains [28, 10, 16, 35]. Various techniques for fine-tuning LLMs have been developed, including the well-known Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) [1, 28]. SFT focuses on tailoring the LLMs’ responses for specific tasks directly with labeled data, whereas RLHF improves LLMs through feedback data reflecting human preferences. In particular, RLHF is pushing the application boundaries of both closed-source [20, 3, 27] and open-source LLMs [28, 33], due to the necessity for polishing the value, fairness, and helpfulness of LLMs in practical scenarios [38, 25, 21].
@@
Existing RLHF methods can be majorly categorized into two classes based on whether the reward signal is explicitly modeled. Reward-based alignment pioneered by OpenAI [21, 1, 28] first trains a reward model from user preferences, typically through Maximum Likelihood Estimation (MLE), and then leverages actor-critic algorithms such as Proximal Policy Optimization (PPO) [24] to tune the SFT model to realize alignment. This approach often requires substantial computational resources and suffers from sample inefficiency [9]. Conversely, another class of methods, known as reward-free alignment, such as Direct Preference Optimization (DPO) [23], Identity Preference Optimization (IPO) [4], and Sequence Likelihood Calibration (SLiC) [36], do not rely on an extra reward model. These approaches offer a more resource-efficient alternative by optimizing the policy directly from preferences, therefore attracting much attention from the academic society. In this work, we commence our analysis with the vanilla DP as a case study, subsequently extending our findings to encompass broader reward-free alignment strategies.
Despite the simplicity and promise of DPO, a variety of phenomena that cannot be clearly understood or explained have been observed and reported in practice. A representative counter-intuitive observation is that the likelihood of both the chosen (i.e., preferred) and the rejected (i.e., less preferred) response will decrease over DPO’s training process [34, 19], while the likelihood of some tokens deviating from the dataset increases [31]. More observations are summarized in Section 2.1. Arguably, purely empirical investigation on applying or improving DPO without a deep exploration and understanding of these phenomena would suffer from low efficiency.
@@
In this paper, we delve into the issues surrounding DPO and its variants from both theoretical and practical perspectives. Our investigation identifies inherent instability in the DPO training process, which we summarize as the 3D-properties: Drastic drop in the likelihood of rejected responses, Degradation into LLM unlearning, and the Dispersion effect on unseen responses. Our analytical framework comprehensively elucidates the reasons behind these phenomena, demonstrating that the 3D-properties significantly undermine DPO’s performance. Moreover, we demonstrate that several DPO variants, such as IPO and SLiC, inherently address these challenges, hence enjoying enhanced performance. Furthermore, our findings confirm that the data distribution of the preference dataset critically influences DPO’s effectiveness. Notably, on-policy DPO exhibits the best performance, which aligns perfectly with recent empirical studies [26, 13].
@@
We propose several regularization methods to enhance the stability of DPO, including adjusting positive and negative weights adaptively and incorporating SFT loss. Our results indicate a potential fundamental trade-off within the DPO algorithm: balancing the mitigation of the 3D-properties against preserving model generalization in preference learning. Furthermore, we thoroughly compare DPO with the state-of-the-art reward-based method RLHF-PPO, revealing that the latter’s superiority may in part stem from its avoidance of 3D-properties. Our experimental approach begins with the design of a toy model to quickly validate our hypotheses, followed by a rigorous test of the actual performance of Baichuan2-13B and Baichuan2-33B [33] on tasks such as mathematical problem solving and instruction following.
@@ 2 Preliminaries
Large Language Model (LLM). An LLM defines a $\theta$-parameterized conditional distribution $\pi_{\theta}(a|x)$, which takes a prompt $x$ as input and produces a response $a$. More specifically, the sampling from LLMs is performed in an auto-regressive manner:
$\pi_{\theta}(a|x)=\prod_{t}\pi_{\theta}(a_{t}|x,a_{1:t-1}),$ (1)
where $a_{t}$ is the $t$-th token in the response $a$ and $a_{1:t-1}$ are tokens in the response before $a_{t}$.
RLHF-PPO. Training LLMs typically involves three stages: Pretraining, SFT, and RLHF. In this section, we outline the standard RLHF-PPO paradigm, widely adopted in advanced research [38, 21].
Beginning with a well-trained SFT model, denoted as $\pi_{0}$, we proceed by sampling two responses from $\pi_{0}$ for each instance in a given prompt set. Subsequently, we compile a set of comparisons $\mathcal{D}=\{(x,a^{+},a^{-})\}$, where $a^{+}$ and $a^{-}$ denote human-preferred and human-dispreferred completions, respectively. The distribution of the preference dataset is assumed to follow the Bradley-Terry model [7], i.e., the probability of response $a^{+}$ is better than $a^{-}$ is given by:
@@
where $\succ$ represents the preference relation, and $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function. To train a reward model $r$, we maximize the log-likelihood of the observed preferences by minimizing the following loss function:
@@
During the RL optimization phase, we update the LM to maximize the return from the learned reward model using the following principle:
@@
where $\pi_{\theta}$ is initialized as $\pi_{0}$ and $\beta$ controls the deviation from the original model. PPO [24] is typically used to solve the problem in practice. Algorithms that optimize the policy using a separate reward model are referred to as reward-based alignment.
@@
DPO. Instead of learning a separate reward model, DPO [23] directly optimizes the policy $\pi_{\theta}$ over preference data. DPO leverages a particular choice of reward model parameterization that enables the extraction of its optimal policy in closed form, without an RL training loop:
$\ell_{\text{DPO}}(\theta)=-\sum_{(x,a^{+},a^{-})}\log\sigma\left[\beta\log\frac{\pi_{\theta}({a}^{+}|x)}{\pi_{0}({a}^{+}|x)}-\beta\log\frac{\pi_{\theta}({a}^{-}|x)}{\pi_{0}({a}^{-}|x)}\right].$ (5)
@@
As shown, DPO leverages logistic regression loss to directly fine-tune the LM on preference data. This approach, along with its various variants [36, 2, 4], is referred to as reward-free alignment due to the elimination of an explicit reward model.
2.1 Underexplored Facts about the DPO Algorithm
@@
Though the absence of the need for additional reward model training makes DPO particularly attractive, there are still many underexplored observations. The most concerning issue is, to the best of our knowledge, few models using DPO (or other reward-free algorithm) have achieved performance comparable to the state-of-the-art closed-source LLMs such as OpenAI’s ChatGPT or Anthropic’s Claude, which reportedly use PPO methods during training. Besides, many other phenomena have been reported but lack comprehensive theoretical explanations. Here we make a summary for clarity.
Fact 1.
@@
During the vanilla DPO training, the likelihood of both the chosen and rejected responses in the preference datasets tends to decrease [19], whereas the likelihood of unseen tokens not appearing in the preference pairs tends to increase [31].
Fact 2.
@@
Compared with RLHF-PPO, the performance of DPO is relatively unstable and sub-optimal [31].
Fact 3.
@@
The performance of DPO is significantly affected by the distribution shift between the model outputs and the preference dataset. In general, on-policy DPO, where both the chosen responses and the rejected response are sampled from the policy model $\pi_{\theta}$, outperforms other variants [26].
In the following section, we will offer theoretical explanations for these facts.
3 The Fundamental Limitation of Vanilla DPO: 3D-Properties
@@
We first identify a critical flaw inherent in the vanilla DPO framework. At first glance, the loss function of standard DPO, as defined in Equation 5, seems to be divided into two segments. The term $\log\frac{\pi_{\theta}(a^{+}|x)}{\pi_{0}(a^{+}|x)}$ aims to increase the likelihood of the chosen response, whereas the term $\log\frac{\pi_{\theta}(a^{-}|x)}{\pi_{0}(a^{-}|x)}$ seeks to reduce the likelihood of the rejected response. Yet, this straightforward interpretation can lead to a serious overlook of core issues, which we define through the 3D-properties of vanilla DPO:
@@
The likelihood of a rejected response changes much more rapidly than that of a chosen response.
Property 2 (Degradation into LLM unlearning).
DPO will gradually fail to control the optimization direction of chosen responses and degrade into solely unlearning the rejected responses as the optimization progresses.
@@
As the training of DPO progresses, the likelihood associated with both chosen and rejected responses will gradually decrease. Concurrently, the likelihood of generating responses not present in the dataset will increase.
In the forthcoming sections, we first provide the theoretical foundations, followed by detailed explanations of the observations discussed in Section 2.1. Subsequently, we introduce a simplified toy model specifically designed to facilitate synthetic experiments, thereby enhancing the persuasiveness of our arguments.
3.1 Theoretical Foundation
The loss function for DPO (Equation 5) can be re-written by:
$\ell^{\text{DPO}}(\theta)=\sum_{(x,a^{+},a^{-})}\log\left(1+\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}\right)^{\beta}\right).$ (6)
@@ 3.1 Theoretical Foundation
$\alpha:=\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\right)^{\beta},\quad\pi^{+}:=\pi(a^{+}|x),\quad\pi^{-}:=\pi(a^{-}|x),\quad z=\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}.$
@@
We have the following observations:
@@
The gradient ratio is:
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}/\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}=\frac{\partial z}{\partial\pi^{-}}/\frac{\partial z}{\partial\pi^{+}}=\frac{\pi^{+}}{\pi^{-}}.$
From the above corollary, we can observe that the gradient with respect to $\partial\pi^{-}$ not only tends to increase but also accelerates in growth as the pair $(\pi^{+},\pi^{-})$ is optimized toward $(1,0)$.
@@
As $\pi^{-}\to 0$ and given that $\beta<1$, $\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}$, which is proportional to $(\pi^{-})^{\beta}$, tends towards zero. Conversely, $\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}$, proportional to $(\pi^{-})^{\beta-1}$, approaches infinity. In this case, the gradient associated with the rejected response becomes exceedingly large, while the gradient for the chosen response diminishes significantly. This dynamic leads to a situation where DPO progressively focuses on unlearning the rejected responses.
@@
When $\pi^{-}$ drastically drops to 00, the gradient on $\pi^{+}$ gradually fails and the likelihood of the chosen response is likely to decrease randomly. As both $\pi^{+}$ and $\pi^{-}$ decrease, the generation probability will randomly disperse into other unseen responses out of the preference dataset.
Based on these theoretical insights, we can further explore the observed facts in Section 2.1. Fact 1 can be explained directly. For Fact 2, we will show that 3D-properties don’t exist in the RLHF-PPO framework in Section 4. For Fact 3, we will demonstrate that the distribution gap between the model’s output and the preference dataset dictates the strength of influence exerted by the 3D-properties. Comparatively, the impact of 3D-properties is less pronounced in on-policy DPO.
3.2 Synthetic Validation in a Toy Model
Figure 1: Toy model setup, preference dataset construction, and results before and after DPO training.
Figure 2: Vanilla DPO performance on toy model.
Before conducting experiments on real LLMs, we introduce a light toy model, which can help us validate Corollary 1 to 3 in a straightforward way.

### v2

@@ Abstract
Aligning large language models (LLMs) with human preferences has gained significant attention, with Proximal Policy Optimization (PPO) as a standard yet computationally expensive method and Direct Preference Optimization (DPO) as a more efficient alternative. While DPO offers simplicity, it remains underutilized in state-of-the-art LLMs, suggesting potential limitations. In this work, we revisit DPO, analyzing its theoretical foundations and empirical performance to bridge this gap. We identify three key properties—termed 3D-properties—that emerge from DPO’s learning process: Drastic drop in rejected response likelihood, Degradation into response suppression, and Dispersion effect on unseen responses. We show that these issues arise from DPO’s optimization dynamics, where the interaction between chosen and rejected response gradients leads to instability. Our findings are supported by experiments on both a controlled toy model and real-world LLM tasks, including mathematical problem-solving and instruction following. To address these challenges, we propose simple regularization techniques that improve training stability and performance. Additionally, we examine how preference data distribution impacts DPO’s effectiveness, offering insights into how alignment models handle out-of-domain (OOD) data. Our work connects these observations to broader research and provides a theoretical explanation for DPO’s limitations. We hope these insights will guide future advancements in reward-model-free preference learning, bringing it closer to reward-model-based approaches.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across a wide range of tasks and domains (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2023; Zhang et al., 2022). Several techniques have been developed for fine-tuning LLMs, most notably Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (Achiam et al., 2023; Touvron et al., 2023). SFT involves directly training LLMs on labeled data to tailor their responses for specific tasks, whereas RLHF refines LLMs by incorporating feedback that aligns their outputs with human preferences. RLHF, in particular, has been instrumental in expanding the application of both closed-source (OpenAI, 2022; Anthropic, 2024; Team et al., 2023) and open-source LLMs (Touvron et al., 2023; Yang et al., 2023), driven by the need to align foundational models with human values and preferences (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022).
@@
Existing RLHF methods can be majorly categorized into two classes based on whether the reward signal is explicitly modeled. Reward-model-based (RM-based) alignment pioneered by OpenAI (Ouyang et al., 2022; Achiam et al., 2023; Touvron et al., 2023) first trains a Reward Model (RM) from user preferences, typically through Maximum Likelihood Estimation (MLE), and then leverages actor-critic algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) to tune the SFT model to realize alignment. This approach often requires substantial computational resources and suffers from sample inefficiency (Choshen et al., 2019). Conversely, another class of methods, known as reward-model-free (RM-free) alignment, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), Identity Preference Optimization (IPO) (Azar et al., 2024), Sequence Likelihood Calibration (SLiC) (Zhao et al., 2023), DPO-positive (Pal et al., 2024) and Simple Preference Optimization (SimPO) (Meng et al., 2024), do not rely on an extra RM. These approaches offer a more resource-efficient alternative by optimizing the policy directly from preferences, therefore attracting much attention from the academic community, where computational resources are often limited.
In this work, we begin our analysis by using the vanilla DPO as a case study, subsequently extending our findings to encompass broader RM-free alignment strategies. Despite its simplicity and promise, DPO has exhibited several perplexing phenomena that remain unclear or underexplained in practice. One notable counter-intuitive observation is that the likelihood of both preferred and rejected responses tends to decrease over the course of DPO training (Yuan et al., 2024; Mitchell, 2023), while the likelihood of certain tokens diverging from the training data increases (Xu et al., 2024a). Additional observations are summarized in Section 2.1. Without a deeper theoretical exploration of these phenomena, purely empirical efforts to apply or improve DPO are likely to face inefficiencies.
@@
Our work identifies the issues surrounding vanilla DPO and its variants from both theoretical and practical perspectives. The analysis reveals inherent instability in the DPO training process, which we encapsulate as the 3D-properties: Drastic drop in the likelihood of rejected responses, Degradation into response suppression, and Dispersion effect on unseen responses. Through our analytical framework, we show that these phenomena stem from the inherent features of DPO’s optimization objective, where the interaction between the gradients of chosen and rejected responses leads to instability and hinders overall performance. Furthermore, our findings confirm that the distribution of preference data critically influences DPO’s effectiveness, with on-policy DPO performing better than off-policy DPO, which is consistent with concurrent empirical studies (Tang et al., 2024; Guo et al., 2024).
@@
To enhance DPO’s stability and performance, we propose several regularization methods, including the adaptive adjustment of weights on the gradients of chosen and rejected responses, as well as incorporating an SFT loss into the objective. Our results suggest a fundamental trade-off within the DPO algorithm: balancing the mitigation of the 3D-properties while preventing LLMs from straying too far from the preference learning paradigm. Additionally, we compare DPO with the state-of-the-art RM-based method, RLHF-PPO, revealing that its superiority stem largely from avoiding the 3D-properties. Our experimental approach begins with the design of a toy model to quickly validate our hypotheses, followed by a rigorous test of the actual performance of real LLMs on tasks such as mathematical problem solving and instruction following.
As this topic has garnered significant attention recently, an increasing number of works are contributing to the discussion. To highlight the contributions of our approach, we compare our findings with several of the most relevant concurrent studies in Section 2.2. A comprehensive review of related works is provided in Appendix A.
@@ 2 Preliminaries
Large Language Model (LLM). An LLM defines a $\theta$-parameterized conditional distribution $\pi_{\theta}(a|x)$, which takes a prompt $x$ as input and produces a response $a$. More specifically, the sampling from LLMs is performed in an auto-regressive manner, $\pi_{\theta}(a|x)=\prod_{t}\pi_{\theta}(a_{t}|x,a_{1:t-1})$, where $a_{t}$ is the $t$-th token in the response $a$ and $a_{1:t-1}$ are tokens in the response before $a_{t}$.
RM-based RLHF. Training LLMs typically involves three stages: Pretraining, SFT, and RLHF. We outline the standard PPO paradigm here, which is a typical RM-based RLHF algorithm. Beginning with a well-trained SFT model, denoted as $\pi_{0}$, we proceed by sampling two responses from $\pi_{0}$ for each instance in a given prompt set. Subsequently, we compile a preferece dataset $\mathcal{D}=\{(x,a^{+},a^{-})\}$, where $a^{+}$ and $a^{-}$ denote human-preferred and human-dispreferred completions, respectively. The distribution of the preference dataset is assumed to follow the Bradley-Terry model (Bradley & Terry, 1952), i.e., the probability of response $a^{+}$ is better than $a^{-}$ is given by:
@@
where $\succ$ represents the preference relation, and $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function. To train a RM $r(\cdot,\cdot)$, we maximize the log-likelihood of the observed preferences by minimizing the following loss function:
@@
During the reinforcement learning phase, we update the LLM to maximize the return from the learned RM using the following objective function:
@@
where $\pi_{\theta}$ is initialized as $\pi_{0}$ and $\beta$ controls the deviation from the original model. PPO (Schulman et al., 2017) is typically used to solve the problem in practice. Algorithms that optimize the policy using a separate RM are referred to as RM-based alignment.
@@
DPO. Instead of learning a separate RM, DPO (Rafailov et al., 2024) directly optimizes the policy $\pi_{\theta}$ over preference data. DPO implicitly leverages a particular choice of RM parameterization that enables the extraction of its optimal policy in closed form, without a reinforcement learning training loop:
$\ell^{\text{DPO}}(\theta)=-\sum_{(x,a^{+},a^{-})}\log\sigma\left[\beta\log\frac{\pi_{\theta}({a}^{+}|x)}{\pi_{0}({a}^{+}|x)}-\beta\log\frac{\pi_{\theta}({a}^{-}|x)}{\pi_{0}({a}^{-}|x)}\right].$ (4)
@@
As shown, DPO leverages logistic regression loss to directly fine-tune the LLM on preference data. This approach, along with its various variants (Zhao et al., 2023; Amini et al., 2024; Azar et al., 2024), is referred to as RM-free alignment due to the elimination of an explicit RM.
2.1 Underexplored Observations about DPO
@@
Though the absence of the need for additional RM training makes DPO particularly attractive, several observations remain underexplored. The most concerning issue is, to the best of our knowledge, few models using DPO (or other RM-free algorithm) have achieved performance comparable to the state-of-the-art closed-source LLMs such as OpenAI’s GPT-4o or Anthropic’s Claude, which reportedly use PPO methods during training. Besides, many other phenomena have been reported but lack comprehensive theoretical explanations. Here we make a summary for clarity.
Observation 1.
@@
During the vanilla DPO training, the likelihood of both the chosen and rejected responses in the preference datasets tends to decrease, whereas the likelihood of unseen tokens not appearing in the preference pairs tends to increase (Mitchell, 2023).
Observation 2.
@@
Compared with RM-based alignment, the performance of DPO is relatively unstable and sub-optimal (Wang et al., 2024).
Observation 3.
@@
The performance of DPO is significantly affected by the distribution shift between the model outputs and the preference dataset. In general, on-policy DPO, where both the chosen responses and the rejected responses are sampled from the policy model $\pi_{\theta}$, outperforms other scenarios (Tang et al., 2024).
2.2 Comparison with Related Contemporary Studies
Several concurrent studies attempt to explain these observations, and we highlight our differences to underscore our contributions. For Observation 1, Feng et al. (2024) shares similar points of view regarding the degradation on gradients but offers limited analysis and lacks experimental validation on real LLMs. In contrast, our work offers a more rigorous and comprehensive analysis supported by theoretical insights and validation across toy models and large-scale real-world LLM experiments. For Observation 2, Xu et al. (2024a) point out that the policy minimizing the PPO loss is a subset of that minimizing the DPO loss, which offers a partial explanation. However, their analysis focuses solely on the endpoint of the optimization and does not examine the dynamic process by which the policy evolves. It leaves unaddressed how unexpected policies emerge during training. In contrast, our gradient analysis offers a comprehensive understanding of the entire optimization trajectory, shedding light on how and why sub-optimal policies arise throughout the DPO training process. For Observation 3, Tang et al. (2024) investigate the performance gap between on-policy and off-policy alignment algorithms from an empirical perspective, while our insights are rooted in theoretical findings.
Our work advances the understanding of these observations, providing critical insights into the underlying mechanisms and reinforcing the findings of these concurrent studies. In the following section, we will present our theoretical explanations for these observations.
3 Fundamental Limitations of Vanilla DPO: 3D-Properties
@@
We first identify a critical flaw inherent in vanilla DPO. At first glance, the loss function of vanilla DPO, as defined in Eq. (4), appears to be composed of two parts: the term $\log\frac{\pi_{\theta}(a^{+}|x)}{\pi_{0}(a^{+}|x)}$ aims to increase the likelihood of chosen responses, while the term $\log\frac{\pi_{\theta}(a^{-}|x)}{\pi_{0}(a^{-}|x)}$ seeks to decrease the likelihood of rejected responses. However, this seemingly straightforward interpretation overlooks significant underlying issues, which we characterize through the 3D-properties of vanilla DPO.
@@
The likelihood of a rejected response tends to shift much more rapidly than that of a chosen response.
Property 2 (Degradation into response suppression).
As optimization progresses, DPO gradually loses its ability to steer the direction of optimizing chosen responses and instead devolves into merely suppressing the rejected responses.
@@
As DPO training progresses, the likelihood of both chosen and rejected responses gradually decreases, while the likelihood of generating out-of-distribution (OOD) responses increases.
These properties are non-trivial, as they reveal inherent challenges in DPO’s optimization process that are not immediately apparent from the loss function alone. These phenomena closely align with the empirical observations we discussed earlier, pointing to the structural limitations of DPO. In the following section, we delve into a theoretical analysis to further explain the origins of these 3D-properties and provide insights into their impact on the optimization trajectory.
3.1 Theoretical Foundations
In this sections, we provide the theoretical foundations for the 3D-properties, followed by detailed explanations of the observations discussed in Section 2.1. The loss function for DPO as shown in Eq. (4) can be re-written by:
$\ell^{\text{DPO}}(\theta)=\sum_{(x,a^{+},a^{-})}\log\left(1+\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\cdot\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}\right)^{\beta}\right).$ (5)
@@ 3.1 Theoretical Foundations
$\alpha:=\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\right)^{\beta},\quad\pi^{+}:=\pi_{\theta}(a^{+}|x),\quad\pi^{-}:=\pi_{\theta}(a^{-}|x),\quad z:=\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}=\frac{\pi^{-}}{\pi^{+}}.$
@@
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}=\frac{\partial\log(1+\alpha z^{\beta})}{\partial z}\frac{\partial z}{\partial\pi^{+}}=\frac{\alpha\beta}{1+\alpha z^{\beta}}z^{\beta-1}\frac{\partial z}{\partial\pi^{+}}=\frac{\alpha\beta}{1+\alpha z^{\beta}}z^{\beta-1}\left[-\frac{\pi^{-}}{(\pi^{+})^{2}}\right],$
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}=\frac{\partial\log(1+\alpha z^{\beta})}{\partial z}\frac{\partial z}{\partial\pi^{-}}=\frac{\alpha\beta}{1+\alpha z^{\beta}}z^{\beta-1}\frac{\partial z}{\partial\pi^{-}}=\frac{\alpha\beta}{1+\alpha z^{\beta}}z^{\beta-1}\left(\frac{1}{\pi^{+}}\right),$
with these simplified forms, we can obtain the following corollaries to explain the properties above:
@@
The ratio of the gradient with respect to the rejected response likelihood $\pi^{-}$ to the gradient with respect to the chosen response likelihood $\pi^{+}$ is equal to the ratio of $\pi^{+}$ to $\pi^{-}$:
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}/\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}=\frac{\partial z}{\partial\pi^{-}}/\frac{\partial z}{\partial\pi^{+}}=-\frac{\pi^{+}}{\pi^{-}},$
which indicates that as $\pi^{+}$ increases and $\pi^{-}$ decreases, the gradient with respect to $\pi^{-}$ grows faster, leading to a more rapid decline in the likelihood of the rejected response.
@@
As $\pi^{-}\to 0$, we have $z\to 0$ and $\frac{\alpha\beta}{1+\alpha z^{\beta}}\to\alpha\beta$. Thus,
$\frac{\partial\ell^{DPO}}{\partial\pi^{+}}\to-\alpha\beta(\pi^{+})^{-\beta-1}(\pi^{-})^{\beta}\to 0,\quad\frac{\partial\ell^{DPO}}{\partial\pi^{-}}\to\alpha\beta(\pi^{+})^{-\beta}(\pi^{-})^{\beta-1}\to\infty,$
given that $\beta<1$ and $\pi^{-}\to 0$. This creates a dynamic where the gradient for the rejected response grows exceedingly large, while the gradient for the chosen response diminishes significantly. As a result, DPO progressively shifts its focus to suppressing the rejected responses and loses the ability to steer the direction of optimizing chosen responses.
@@
When $\pi^{-}$ drastically drops to 00, the gradient on $\pi^{+}$ fails and the likelihood of the chosen response is likely to decrease along with the rejected response as they often share many similar tokens and patterns. The constancy of the sum of probabilities implies that as both $\pi^{+}$ and $\pi^{-}$ decrease, the likelihood will randomly disperse into other unseen responses out of the preference dataset.
Based on these theoretical insights, we can further explore and explain the observations discussed in Section 2.1. Observation 1 is directly explained by Corollary 3. For Observation 2, we will show that 3D-properties do not manifest during the RM training process in the RM-based alignment pipeline. Regarding Observation 3, we will demonstrate that the distribution gap between the LLM’s original outputs and the preference dataset plays a crucial role in determining the influence of the 3D-properties. The impact of these properties is notably less pronounced in on-policy DPO, where the preference dataset is sampled directly from the policy model’s outputs. In the following sections, we will delve deeper into each of these statements and provide empirical validation.
3.2 Synthetic Validation with a Toy Model
Figure 1: Toy model setup. Top left: the optimal policy where the highlighted blocks represent optimal responses. Top right: preference dataset construction. Lower left: the initialization of the SFT model. Lower right: policy output after DPO training.
Figure 2: Dynamic optimization process with vanilla DPO using the toy model. Left: likelihood dynamics over training epochs. The blue curve represents the average likelihood of chosen responses, yellow shows the minimum for chosen responses, green represents the average for rejected responses, red shows the maximum for rejected responses, and purple represents the average for unseen responses. Middle: dynamics of averaged $\frac{\partial\ell^{DPO}}{\partial\pi^{+}}$ and $\frac{\partial\ell^{DPO}}{\partial\pi^{-}}$ over training epochs. Right: likelihood dynamics over training epochs on a log scale, highlighting the drastic drop in the likelihood of rejected responses.
In this section, we introduce a simplified toy model specifically designed to facilitate synthetic experiments, thereby enhancing the persuasiveness of our arguments from Corollary 1 to 3. Then we conduct experiments on real LLMs.